OpenStudy (anonymous):

What are the x- and y-intercepts of the graph of y = x2 + x - 6? (1 points) Answer x-intercepts: (3, 0) and (2, 0) y-intercept: (0, 6) x-intercepts: (3, 0) and (-2, 0) y-intercept: (0, -6) x-intercepts: (-3, 0) and (2, 0) y-intercept: (0, -6) x-intercepts: (-3, 0) and (2, 0) y-intercept: (0, 6)

OpenStudy (anonymous):

to find the y intercept set x=0 on y = x2 + x - 6, what is y=__? to find the x intercept set y=0 on y = x2 + x - 6=0, what is x=__?hint factor it or use the quadratic formula... :D

OpenStudy (anonymous):

a) to find the y intercept set x=0 on y = x2 + x - 6, what is y=__? solution: set x=0 y = x2 + x - 6 y = (0)^2 + (0 - 6 =___?

OpenStudy (anonymous):

b) to find the x intercept set y=0 on y = x2 + x - 6=0, what is x=__?hint factor it or use the quadratic formula... :D sol: set y=0 y = x2 + x - 6=0 try factoring (x +__?)(x -__?)=0 or maybe use the quadratic equation.. :D

OpenStudy (anonymous):

I got C @mark_o.

OpenStudy (anonymous):

yeheyyyyy correct :D

OpenStudy (anonymous):

thank yoU!!!!!!!!!!!!!! :)

OpenStudy (anonymous):

you are correct :D (x +3)(x -2)=0 therefore x=-3, 2

OpenStudy (anonymous):

YW and good luck now... :D have fun :D